### 30 Dec 2023 v1

@@ Abstract
The sparsity of reward feedback remains a challenging problem in online deep reinforcement learning (DRL). Previous approaches have utilized temporal credit assignment (CA) to achieve impressive results in multiple hard tasks. However, many CA methods relied on complex architectures or introduced sensitive hyperparameters to estimate the impact of state-action pairs. Meanwhile, the premise of the feasibility of CA methods is to obtain trajectories with sparse rewards, which can be troublesome in sparse-reward environments with large state spaces. To tackle these problems, we propose a simple and efficient algorithm called Policy Optimization with Smooth Guidance (POSG) that leverages a small set of sparse-reward demonstrations to make reliable and effective long-term credit assignments while efficiently facilitating exploration. The key idea is that the relative impact of state-action pairs can be indirectly estimated using offline demonstrations rather than directly leveraging the sparse reward trajectories generated by the agent. Specifically, we first obtain the trajectory importance by considering both the trajectory-level distance to demonstrations and the returns of the relevant trajectories. Then, the guidance reward is calculated for each state-action pair by smoothly averaging the importance of the trajectories through it, merging the demonstration’s distribution and reward information. We theoretically analyze the performance improvement bound caused by smooth guidance rewards and derive a new worst-case lower bound on the performance improvement. A comprehensive evaluation of POSG is conducted by benchmarking it with several state-of-the-art RL methods in four different sparse-reward environments with discrete and continuous action spaces. Extensive results demonstrate POSG’s significant advantages in control performance and convergence speed compared to benchmark DRL algorithms. Notably, the specific metrics and quantifiable results are investigated to demonstrate the superiority of POSG.
@@
deep reinforcement learning, credit assignment, sparse rewards, offline demonstrations, policy optimization
@@ 1 Introduction
In recent years, deep reinforcement learning (RL) has achieved impressive results in solving sequential decision-making tasks in various domains, such as the Arcade Learning Environment Mnih2015HumanlevelCT ; bellemare2013arcade ; ecoffet2021first , the game of Go Silver2016MasteringTG , locomotive continuous control Lillicrap2016ContinuousCW ; Schulman2015TrustRP ; fujimoto2018addressing , and robotic navigation Florensa2017StochasticNN ; zhanghierarchical ; li2022active . Despite these famous success stories, reinforcement learning remains a challenging problem in sparse or delayed reward settings yang2021exploration ; gangwani2020learning because it is intractable to maintain an acceptable trade-off of exploration against exploitation in such tasks with sparse or delayed rewards guo2020memory . Typical DRL algorithms suffer from the exploration issue because fully exploring the entire state-action space remains difficult and cannot be guaranteed spontaneously, especially when environmental rewards are sparse jing2020reinforcement .
@@
Many research advances have assisted policy exploration of agents using expert demonstrations zhou2020watch . An intuitive RLfD approach enhances RL by data augmentations, which maintains the expert demonstrations in a replay buffer for value estimation hester2018deep ; vecerik2017leveraging ; gulcehre2019making ; libardi2021guided . Some RLfD methods utilize demonstrations to pre-train the policy by supervised learning silver2016mastering ; cruz2017pre . These algorithms force the agent’s policy to conform to the expert’s policy and do not reuse them during the policy optimization procedure. Recent RLfD studies draw inspiration from imitation learning and encourage the agent to mimic the demonstrated actions kang2018policy ; Sun2018TruncatedHP ; jing2020reinforcement ; rengarajan2022reinforcement . Specifically, these methods either augment the original RL loss function with a divergence regularization term or design a new shaping reward derived from a distribution divergence function to force expert-alike exploration.
Whereas most existing RLfD methods cannot simultaneously exploit the demonstrations’ distributional information and reward signals. More specifically, data augmentations with demonstrations only exploit the reward information in the expert demonstrations to learn the $Q$ function and indirectly encourage the agent to perform the demonstrated behaviors. These methods ignore demonstration distribution information, causing a potential information bottleneck jing2020reinforcement ; rengarajan2022reinforcement . Furthermore, these approaches require reward signals to be dense. On the other hand, if the distributional information is only considered kang2018policy ; jing2020reinforcement ; rengarajan2022reinforcement , we require a large number of expert demonstrations to estimate the expert distribution. Due to this drawback, the RLfD demonstration data efficiency is often low, and RLfD places a heavy burden on the demonstration quality, limiting the realistic application of RLfD methods.
This study designs a simple and feasible reinforcement learning method, called Policy Optimization with Smooth Guidance rewards (POSG), which can be easily integrated into the existing RL algorithms. POSG enables reliable long-term credit assignments (CA) and expands the application scope of CA approaches since our method does not depend on complex architectures and does not require obtaining sparse-reward trajectories before credit assignment. Meanwhile, POSG overcomes the drawbacks of RLfD methods and reduces RLfD demonstration requirements. The key insight is that leveraging offline demonstration trajectories measures the impact of current state-action pairs, which merges demonstrations’ distribution information with their reward signals. More specifically, the trajectory importance is computed based on the maximum mean discrepancy (MMD) distance to demonstrations and the returns of associated trajectories. We then obtain the guidance reward for each state-action pair with a smooth weighted average of the trajectory importance. Furthermore, we deduce a new worst-case lower bound for POSG policy optimization and provide a theoretical performance improvement guarantee. Extensive experimental results indicate that POSG outperforms other baseline algorithms in different environments.
@@
    To avoid the CA’s dilemma, we propose a simple, effective, and efficient RL method that utilizes offline sparse-reward demonstrations to achieve reliable credit assignments in sparse-reward settings.
@@
    POSG is capable of using a single demonstration trajectory to solve the exploration problem of RL and improve the sample efficiency of RL.
@@
    A theoretical analysis demonstrates the effectiveness of information fusion, and a new worst-case lower bound is deduced to provide a performance improvement guarantee for POSG.
@@
    POSG outperforms other state-of-the-art RL algorithms in different environments. Notably, the specific metrics and quantifiable results are investigated to demonstrate the superiority of POSG.
@@ 2 Related Work
Reinforcement Learning from Demonstrations. RLfD combines RL with expert demonstration data to improve policy exploration and accelerate learning. To accomplish this goal, RLfD trains a behavioral policy with a state-action visitation distribution similar to the demonstrator’s. Early RLfD algorithms sample replay buffers containing expert demonstrations and self-generated data simultaneously to enhance the learning ability of the agent hester2018deep ; vecerik2017leveraging . Self-imitation learning (SIL) methods oh2018self ; gangwani2019learning ; libardi2021guided train the agent to imitate its own past experiences only when the return of the previous episode is greater than the value estimate of the agent or returns of trajectories in the replay buffer. Episodic reinforcement learning methods utilize episodic memories to estimate the value of the state precisely and propagate its value to the previous states pritzel2017neural ; blundell2016model ; lin2018episodic ; lee2019sample ; hu2021generalizable ; le2021model . Many other previous works introduce novel methods that enable the agent learns a range of diverse exploratory policies based on episodic memory guo2020memory ; badia2020never . This study uses demonstration experiences in the state-action space to construct a shaping function, which can easily be integrated with existing RL methods, such as PPO schulman2017proximal .
Imitation Learning (IL). Imitation learning aims to learn a control policy that imitates the behaviors of experts and outputs the same action when the agent receives the observation that occurred in the demonstration data set. Behavior Clone (BC) kumar2021should is viewed as a simple IL approach that learns a control policy from demonstration data by supervised learning. However, the BC method usually suffers from the problem of distribution shift ross2011reduction . In Inverse Reinforcement Learning (IRL), a reward function is first learned from expert demonstrations, and then it will be used to solve forward RL problems Ng2000AlgorithmsFI ; ziebart2008maximum . Meanwhile, Generative Adversarial Imitation Learning (GAIL) Ho2016GenerativeAI formulates the IL problem as a distribution matching task, which is combined with Generative Adversarial Networks (GAN) goodfellow2020generative . However, these IL methods depend on the availability of sufficient and high-quality expert demonstrations. Instead, our method mitigates this serious problem by combining RL with demonstration data.
@@ 3.1 Reinforcement Learning
A typical RL problem is modeled as an infinite-horizon Markov decision process with discrete time, which can be defined as a tuple $M=(\mathcal{S},\mathcal{A},P,r,\rho_{0},\gamma)$. Here, $\mathcal{S}$ is a discrete or continuous state space, $\mathcal{A}$ denotes a discrete (or continuous) action space, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Pi(\mathcal{S})$ is the transition probability distribution, where $\Pi(\mathcal{S})$ is the space of probability distributions over the state $\mathcal{S}$. In addition, $r_{e}:\mathcal{S}\times\mathcal{A}\rightarrow[R_{min},R_{max}]$ is the reward function, in which we assume that the minimum and maximum value of the reward function is $R_{min}$ and $R_{max}$, respectively. Furthermore, $\rho_{0}$ is the initial state distribution, and $\gamma\in[0,1]$ is a discount factor. A stochastic policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ parameterized by $\theta$, maps the state space $\mathcal{S}$ to the set of probability distributions over the action space $\mathcal{A}$. Generally, the optimization objective of RL is to find a policy $\pi_{\theta}$ that maximizes the expected discounted return:
@@
$Q_{\pi}(s_{t},a_{t})=\mathbb{E}_{s_{t+1},a_{t+1}}\left[\sum_{l}^{\infty}\gamma^{l}r_{e}(s_{t+l},a_{t+l})\right],$ (2)
@@
$V_{\pi}(s_{t})=\mathbb{E}_{a_{t},s_{t+1},a_{t+1}}\left[\sum_{l}^{\infty}\gamma^{l}r_{e}(s_{t+l},a_{t+l})\right],$ (3)
where $s_{0}\sim\rho_{0}(s_{0})$, $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$, and $s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})$. Then, the advantage function is expressed as:
@@ 3.2 Maximum Mean Discrepancy
In this paper, we treat trajectories as deterministic policy distributions and use the maximum mean discrepancy (MMD) to measure the difference (or similarity) between trajectories empirically gretton2006kernel ; Gretton2012OptimalKC ; masood2019diversity . Given any two trajectories $\tau$ and $\upsilon$, we can use the maximum mean discrepancy to measure the distance between them:
@@
Formally, the function $g$ gives us the feasibility that adapts the focus of the ${\rm MMD}$ metric for different aspects, such as state visits, action choices, or both, to different downstream tasks. In our experiments, we define the distance from trajectory to the demonstration dataset $D(\tau,\mathcal{M}_{E})$ as follows:
$D(\tau,\mathcal{M}_{E})=\min_{\tau_{E}\in\mathcal{M}_{E}}\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})$ (7)
@@ 4 Proposed Approach
In this section, we present Policy Optimization with Smooth Guidance rewards (POSG), a computation-efficient framework for achieving reliable long-term credit assignments. The key insight is that we propose to exploit sparse-reward demonstration trajectories to measure the relative importance of current state-action pairs. Our approach merges the distributional information of demonstrations with the return information of trajectories.
@@ 4.1 Smooth Guidance Reward Design
This section defines the trajectory importance based on a novel trajectory-level distance measurement and the trajectory return. In this manner, this importance integrates the distributional information of demonstrations with reward signals of trajectories. We can then fuse these two pieces of information into the guidance reward by a smooth weighted average of the trajectory importance.
We first introduce a probability density model for trajectories through a specific state-action pair to achieve this goal. Let $\pi(a|s)$ denote a behavioral policy and the trajectory distribution $p_{\pi}$ induced by $\pi$ be expressed as:
@@
Given a state-action pair $(s,a)$, we can sample trajectories containing $(s,a)$ or a relevant information subset of $(s,a)$ according to the distribution $p_{\pi}(\tau)$. For clarity, we refer to $(s,a)$ or the relevant information subset of $(s,a)$ as $x$ in the following text. In other words, suppose $x=f(s,a)$, then $x=(s,a)$ when $f$ is the identified map, and $x\neq(s,a)$ when $f$ is a general feature extraction function and not the identified map. We design the following probability density model for trajectories including $x$:
$p_{\pi}(\tau|x):=\frac{p_{\pi}(\tau)\mathbb{I}[x\in f(\tau)]}{\int_{\tau}p_{\pi}(\tau)\mathbb{I}[x\in f(\tau)]\mathrm{d}\tau},$ (9)
where $f(\tau):=(f[(s_{0},a_{0})],f[(s_{1},a_{1})],\cdots)$, and $\mathbb{I}$ is the indicator function. In this manner, $p_{\pi}(\tau|x)$ is a conditional probability density function of trajectories through $x$.
@@
$\omega(\tau_{0}|\mathcal{M}_{E}):=\frac{e^{-kd}}{\int_{\tau}p_{\pi}(\tau)e^{-kd}\mathrm{d}\tau},$ (10)
where $d=D(\tau_{0},\mathcal{M}_{E})$ and $k$ is a predefined positive constant. It is worth noting that $D(\tau_{0},\mathcal{M}_{E})$ is defined in Section 3.2. Similar to $p_{\pi}(\cdot)$, $\omega(\cdot|\mathcal{M}_{E})$ can be considered as a conditional probability density function over the trajectory space. $\omega(\cdot|\mathcal{M}_{E})$ satisfies the conditions of probability distribution $\omega_{\pi}(\tau|\mathcal{M}_{E})\geq 0,\forall\tau$ and $\int_{\tau}p_{\pi}(\tau)\omega(\tau|\mathcal{M}_{E})\mathrm{d}\tau=1$. According to $\omega_{\pi}(\tau|\mathcal{M}_{E})$, a trajectory closer to $\mathcal{M}_{E}$ is granted more weight by $\omega_{\pi}(\cdot|\mathcal{M}_{E})$. Based on the trajectory weight function $\omega_{\pi}(\tau|\mathcal{M}_{E})$, we can define the trajectory importance as follows:
@@
Here, $\tau_{E}=\operatorname{arg\,min}_{\tau_{E}\in\mathcal{M}_{E}}\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})$ is the demonstration trajectory in $\mathcal{M}_{E}$ closest to $\tau$ under the MMD distance metric, and $R(\cdot)$ denotes the return value of the trajectory. The two constants $\alpha$ and $\beta$ in Eq. (12) are non-negative, and their linear sum is $1$, i.e., satisfies the condition $\alpha+\beta=1$. In this manner, the trajectory importance integrates the distribution information of the demonstrations with the returns of the trajectories. Finally, we can define the smooth guidance rewards as follows:
$r_{i}(s_{t},a_{t})=\int_{\tau}p_{\pi}(\tau|x_{t})I(\tau)\mathrm{d}\tau,$ (13)
The guidance reward $r_{i}(s_{t},a_{t})$ allocated to each state-action pair is the expectation of the product of $R_{j}(\tau)$ and $\omega(\tau|\mathcal{M}_{E})$ under the conditional distribution $p_{\pi}(\tau|x)$. The joint return $R_{j}(\tau)$ is the weighted average sum of the environmental returns of $\tau$ and $\tau_{E}$. Moreover, we use $\omega(\tau|\mathcal{M}_{E})$ to adjust its contribution to the guidance reward $r_{i}(s_{t},a_{t})$. Specifically, the closer the trajectory is from $\mathcal{M}_{E}$, the larger the weight of $\tau$. Therefore, such a trajectory will contribute more to the smooth guidance reward. A new smoothed RL objective is obtained by inserting Eq. (9) in the standard RL objective:
@@
Given a trajectory with the outcome, how can we determine the relevance of each state-action pair toward achieving this outcome in tasks with sparse or delayed rewards? This is the central topic that temporal credit assignment copes with. Our method can potentially perform long-term credit assignments. Inspired by RLfD methods, we propose to realize this goal by using offline demonstration data in sparse reward settings. Instead of learning complex models to reveal the relevance of action to the future result or adopting contribution analysis methods to achieve return redistribution, our method learns a guidance reward function to assist policy optimization.
Our method can be regarded as a simple redistribution method - the guidance reward for each state-action pair is obtained by calculating the weighted average of the trajectory returns based on the MMD distance between trajectories. In particular, this design of smooth guidance rewards considers the distributional information of demonstrations and the trajectory returns simultaneously. More importantly, our method does not require obtaining trajectories with sparse rewards before credit, which expands the application scope of CA approaches and provides a new idea to CA.

### 10 Apr 2024 v2

@@ Abstract
The sparsity of reward feedback remains a challenging problem in online deep reinforcement learning (DRL). Previous approaches have utilized offline demonstrations to achieve impressive results in multiple hard tasks. However, these approaches place high demands on demonstration quality, and obtaining expert-like actions is often costly and unrealistic. To tackle these problems, we propose a simple and efficient algorithm called Policy Optimization with Smooth Guidance (POSG), which leverages a small set of state-only demonstrations (where only state information is included in demonstrations) to indirectly make approximate and feasible long-term credit assignments and facilitate exploration. Specifically, we first design a trajectory-importance evaluation mechanism to determine the quality of the current trajectory against demonstrations. Then, we introduce a guidance reward computation technology based on trajectory importance to measure the impact of each state-action pair. We theoretically analyze the performance improvement caused by smooth guidance rewards and derive a new worst-case lower bound on the performance improvement. Extensive results demonstrate POSG’s significant advantages in control performance and convergence speed in four sparse-reward environments, including the grid-world maze, Hopper-v4, HalfCheetah-v4, and Ant maze. Notably, the specific metrics and quantifiable results are investigated to demonstrate the superiority of POSG.
@@
deep reinforcement learning, sparse rewards, state-only demonstrations, policy optimization
@@ 1 Introduction
In recent years, deep reinforcement learning (RL) has demonstrated remarkable accomplishments in tackling sequential decision-making challenges across diverse domains, including the Arcade Learning Environment Mnih2015HumanlevelCT ; bellemare2013arcade ; ecoffet2021first , the game of Go Silver2016MasteringTG , continuous locomotive control Lillicrap2016ContinuousCW ; Schulman2015TrustRP ; fujimoto2018addressing , and robotic navigation Florensa2017StochasticNN ; zhanghierarchical ; li2022active . Despite these celebrated achievements, reinforcement learning remains a formidable task in scenarios characterized by sparse or delayed rewards yang2021exploration ; gangwani2020learning , primarily due to the challenge of striking a balance between exploration and exploitation in environments with sparse or delayed rewards guo2020memory . Conventional deep RL algorithms confront exploration difficulties, as fully exploring the entire state-action space is often unfeasible and cannot be spontaneously guaranteed, particularly in settings with sparse environmental rewards jing2020reinforcement .
@@
Many research advances have assisted policy exploration of agents by learning from demonstrations (LfD) zhou2020watch . An intuitive LfD approach enhances RL by data augmentations, which maintains the expert demonstrations in a replay buffer for value estimation hester2018deep ; vecerik2017leveraging ; gulcehre2019making ; libardi2021guided . Some LfD methods utilize demonstrations to pre-train the policy by supervised learning silver2016mastering ; cruz2017pre . These algorithms force the agent’s policy to conform to the expert’s policy and do not reuse them during the policy optimization procedure. Recent LfD studies draw inspiration from imitation learning and encourage the agent to mimic the demonstrated actions kang2018policy ; Sun2018TruncatedHP ; jing2020reinforcement ; rengarajan2022reinforcement . Specifically, these methods either augment the original RL loss function with a divergence regularization term or design a new shaping reward derived from a distribution divergence function to force expert-alike exploration.
In summary, the limitations of both CA and LfD approaches impede their practicality in sparse-reward settings. Firstly, it can be troublesome for CA methods to obtain highly rewarded trajectories in sparse-reward environments with large state spaces. Moreover, CA methods pose a computational burden and face challenges in accurately estimating the influence of individual state-action pairs. Secondly, most LfD methods demand high-quality samples and rely on flawless and sufficient demonstrations, including complete state-action information. Our study lies at the intersection of CA and LfD methods, aiming to mitigate their shortcomings. This paper introduces a straightforward and practical reinforcement learning approach named Policy Optimization with Smooth Guidance (POSG), designed for seamless integration into existing RL algorithms. Our method facilitates approximate long-term credit assignments and extends the applicability of CA approaches. Furthermore, this method addresses the limitations of LfD approaches by introducing a dense demonstration-based guidance reward function, thereby reducing the requirements for LfD demonstrations.
The proposed approach leverages only a few, or even a single state-only demonstration trajectory, to indirectly estimate the impact of each state-action pair. The fundamental insight lies in utilizing state-only demonstration trajectories to assess the impact of current state-action pairs, thereby amalgamating state distribution information from demonstrations with their associated reward signals. We introduce two technologies to compute the impact of each state-action pair: a trajectory-importance evaluation mechanism and a guidance reward computation technology. Specifically, the trajectory-importance evaluation mechanism estimates trajectory importance based on the maximum mean discrepancy (MMD) distance to demonstrations and the returns of corresponding trajectories. Subsequently, the guidance reward for each state-action pair is derived through a smooth weighted average of trajectory importance. Furthermore, we establish a novel worst-case lower bound for POSG policy optimization and offer a theoretical guarantee of performance improvement. Extensive experimental evaluations demonstrate that POSG surpasses other baseline algorithms across a discrete grid-world maze and three continuous locomotion control tasks.
@@
    To avoid the CA’s dilemma, we propose an RL method that utilizes offline state-only demonstrations to achieve approximate credit assignments in sparse-reward settings.
@@
    POSG can use a single state-only demonstration trajectory to solve RL’s exploration problem and improve RL’s sample efficiency, demonstrated experimentally in Section 7.4.1.
@@
    A new worst-case lower bound is deduced to provide a performance improvement guarantee for POSG.
@@
    This study discusses the superior performance of POSG over other state-of-the-art RL algorithms across a discrete grid-world maze and three continuous locomotion control tasks. Significantly, the analysis focuses on specific metrics and quantifiable results to illustrate the superiority of POSG.
@@ 2 Related Work
Learning from Demonstrations. LfD combines RL with expert demonstration data to improve policy exploration and accelerate learning. LfD trains a behavioral policy with a state-action visitation distribution similar to the demonstrator’s to accomplish this goal. Early LfD algorithms sample replay buffers containing expert demonstrations and self-generated data simultaneously to enhance the learning ability of the agent hester2018deep ; vecerik2017leveraging . Self-imitation learning (SIL) methods oh2018self ; gangwani2019learning ; libardi2021guided train the agent to imitate its own past experiences only when the return of the previous episode is greater than the value estimate of the agent or returns of trajectories in the replay buffer. Episodic reinforcement learning methods utilize episodic memories to estimate the value of the state precisely and propagate its value to the previous states pritzel2017neural ; blundell2016model ; lin2018episodic ; lee2019sample ; hu2021generalizable ; le2021model . Many other previous works introduce novel methods that enable the agent learns a range of diverse exploratory policies based on episodic memory guo2020memory ; badia2020never . This study uses demonstration experiences in the state-action space to construct a shaping function, which can easily be integrated with existing RL methods, such as PPO schulman2017proximal .
State-Only Imitation Learning (IL). Imitation learning aims to learn a control policy that imitates the behaviors of experts and outputs the same action when the agent receives the observation that occurred in the demonstration data set. The state-only IL method is a branch of imitation learning, where the requirement for the demonstrator’s action information is alleviated. This approach expands the scope of the realistic application of IL. GAIfO torabi2018generative proposes to learn a policy to output actions that lead to similar effects as demonstrations. I2L Gangwani2020StateonlyIW and SAIL Liu2019StateAI train the agent by minimizing the distance between state visitation distributions of the current policy and demonstrations. AILO Gangwani2022ImitationLF leverages demonstrations only containing observations to train an intermediary policy whose state transitions are close to the expert dataset.
Metrics to Compute the Distance between Policies or Trajectories. Many studies propose to compute the distance between policies using Kullback–Leibler (KL), Bregman, or $f$-divergence lee2019efficient ; ghasemipour2020divergence ; wang2019divergence . Rényi divergence is introduced to compute the discrepancy of state visitation between different trajectories Yuan2022RewardingEV ; however, this divergence is a parametric distance measure, and this approach employs $k$-NN estimator to efficiently estimate the parameters of distributions. MADE maximizes the derivation of state-action visitations of the current policy from the explored regions of previous trajectories zhang2021made . However, MADE relies on the estimation of the state-action visitation density, which can be non-trivial in high-dimensional control tasks. Similar to our method, CQL kumar2020conservative and MCPO jing2020reinforcement adopt the MMD metric as the discrepancy measure; however, these algorithms require both state and action information of demonstrations to update the policy parameter.
@@ 3.1 Reinforcement Learning
A typical RL problem is modeled as an infinite-horizon Markov decision process with discrete time, which can be defined as a tuple $M=(\mathcal{S},\mathcal{A},P,r,\rho_{0},\gamma)$. Here, $\mathcal{S}$ is a discrete or continuous state space, $\mathcal{A}$ denotes a discrete (or continuous) action space, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Pi(\mathcal{S})$ is the transition probability distribution, where $\Pi(\mathcal{S})$ is the space of probability distributions over the state $\mathcal{S}$. In addition, $r_{e}:\mathcal{S}\times\mathcal{A}\rightarrow[R_{min},R_{max}]$ is the environmental reward function, in which we assume that the minimum and maximum value of the reward function is $R_{min}$ and $R_{max}$, respectively. Furthermore, $\rho_{0}$ is the initial state distribution, and $\gamma\in[0,1]$ is a discount factor. A stochastic policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ parameterized by $\theta$, maps the state space $\mathcal{S}$ to the set of probability distributions over the action space $\mathcal{A}$. Generally, the optimization objective of RL is to find a policy $\pi_{\theta}$ that maximizes the expected discounted return:
@@
$Q_{\pi}(s_{t},a_{t})=\mathbb{E}_{s_{t+1},a_{t+1},\cdots}\left[\sum_{l=0}^{\infty}\gamma^{l}r_{e}(s_{t+l},a_{t+l})\right],$ (2)
@@
$V_{\pi}(s_{t})=\mathbb{E}_{a_{t},s_{t+1},a_{t+1},\cdots}\left[\sum_{l=0}^{\infty}\gamma^{l}r_{e}(s_{t+l},a_{t+l})\right],$ (3)
where $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$, and $s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})$. Then, the advantage function is expressed as:
@@ 3.2 Maximum Mean Discrepancy
In this paper, we treat trajectories as deterministic policy distributions and use the maximum mean discrepancy (MMD) to measure the difference (or similarity) between trajectories empirically gretton2006kernel ; Gretton2012OptimalKC ; masood2019diversity . MMD is a non-parametric test statistic, and different from Kullback-Leibler (KL), Jensen–Shannon (JS), or $f$-divergences, calculating the MMD metric does not require the distribution parameters. This property makes it suitable for the problem considered in this study where trajectories are treated as deterministic policies.
Given any two trajectories $\tau$ and $\upsilon$, we can use the maximum mean discrepancy to measure the distance between them:
@@
Formally, the function $g$ gives us the feasibility that adapts the focus of the ${\rm MMD}$ metric for different aspects, such as state visits, action choices, or both, to different downstream tasks. In our experiments, we define the distance $D(\tau,\mathcal{M}_{E})$ from trajectory to the demonstration dataset $\mathcal{M}_{E}$ as follows:
$D(\tau,\mathcal{M}_{E})=\min_{\tau_{E}\in\mathcal{M}_{E}}\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H}).$ (7)
@@ 4 Proposed Approach
This section presents Policy Optimization with Smooth Guidance (POSG), a computation-efficient framework for achieving approximate long-term credit assignments. The key insight is that we propose to exploit state-only demonstration trajectories to measure the relative importance of current state-action pairs. Our approach merges the distributional information of demonstrations with the return information of trajectories.
@@ 4.1 Smooth Guidance Reward Design
This section defines the trajectory importance based on a novel trajectory-level distance measurement and the trajectory return. In this manner, this importance integrates the state distribution information of demonstrations with reward signals of trajectories. We can then fuse these two pieces of information into the guidance reward by a smooth weighted average of the trajectory importance.
To achieve this goal, we first introduce a probability density model for trajectories through a specific state-action pair. Let $\pi(a|s)$ denote a behavioral policy and the trajectory distribution $p_{\pi}$ induced by $\pi$ be expressed as:
@@
Given a state’s information subset $x$, we can sample trajectories containing $x$ according to the distribution $p_{\pi}(\tau)$. Mathematically, suppose $x=f(s)$, then $x=s$ when $f$ is the identified map, and $x\neq s$ when $f$ is a general feature extraction function and not the identified map. In our experiments, $x$ is the coordinate $c$ of the center of mass (CoM). We design the following probability density model for trajectories including $x$:
$p_{\pi}(\tau|x):=\frac{p_{\pi}(\tau)\mathbb{I}[x\in f(\tau)]}{\int_{\upsilon}p_{\pi}(\upsilon)\mathbb{I}[x\in f(\upsilon)]\mathrm{d}\upsilon},$ (9)
where $f(\tau):=\{f(s_{0}),f(s_{1}),\cdots\}$, $\mathbb{I}$ is the indicator function. $\tau$ is the current target trajectory and $\upsilon$ is an arbitrary trajectory, all sampled from $p_{\pi}$. In this manner, $p_{\pi}(\tau|x)$ is a conditional probability density function of trajectories through $x$ and describes the probability of $\tau$ among all trajectories through $x$.
@@
$\omega(\tau|\mathcal{M}_{E}):=\frac{e^{-kd(\tau)}}{\int_{\upsilon}p_{\pi}(\upsilon)e^{-kd(\upsilon)}\mathrm{d}\upsilon},$ (10)
where $d(\cdot)=D(\cdot,\mathcal{M}_{E})$, and $\tau$ and $\upsilon$ have the same meanings as those in Eq. (9). $k$ is a predefined positive constant to adjust the value of $\omega$. Specifically, by choosing the suitable value of $k$, we can smooth the value of $\omega$ and prevent its value corresponding to a certain trajectory from being too large. It is worth noting that $D(\tau,\mathcal{M}_{E})$ is defined in Section 3.2. Similar to $p_{\pi}(\cdot)$, $\omega(\cdot|\mathcal{M}_{E})$ can be considered as a conditional probability density function over the trajectory space. $\omega(\cdot|\mathcal{M}_{E})$ satisfies the conditions of probability distribution $\omega_{\pi}(\tau|\mathcal{M}_{E})\geq 0,\forall\tau$ and $\int_{\tau}p_{\pi}(\tau)\omega(\tau|\mathcal{M}_{E})\mathrm{d}\tau=1$. According to $\omega_{\pi}(\tau|\mathcal{M}_{E})$, a trajectory closer to $\mathcal{M}_{E}$ is granted more weight by $\omega_{\pi}(\cdot|\mathcal{M}_{E})$. Based on the trajectory weight function $\omega_{\pi}(\tau|\mathcal{M}_{E})$, we can define the trajectory importance as follows:
@@
Here, $R(\cdot)$ denotes the return value of the trajectory. $\tau_{E}=\operatorname{arg\,min}_{\tau_{E}\in\mathcal{M}_{E}}\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})$ is the state-only demonstration in $\mathcal{M}_{E}$ closest to $\tau$ under the MMD distance metric, and $\mathcal{H}$ a reproducing kernel Hilbert space as described in Section 3. The two constants $\alpha$ and $\beta$ in Eq. (12) are non-negative, and their linear sum is $1$, i.e., satisfies the condition $\alpha+\beta=1$. In this manner, the trajectory importance integrates the state distribution information of the demonstrations with the returns of the trajectories. Finally, we can define the smooth guidance rewards for a state-action pair $(s,a)$ as follows:
$r_{i}(s,a)=\int_{\tau}p_{\pi}(\tau|x)I(\tau)\mathrm{d}\tau.$ (13)
Note that $x=f(s)$. The guidance reward $r_{i}(s,a)$ allocated to each state-action pair is the expectation of the product of $R_{j}(\tau)$ and $\omega(\tau|\mathcal{M}_{E})$ under the conditional distribution $p_{\pi}(\tau|x)$. The joint return $R_{j}(\tau)$ is the weighted average sum of the environmental returns of $\tau$ and $\tau_{E}$. Moreover, we use $\omega(\tau|\mathcal{M}_{E})$ to adjust its contribution to the guidance reward $r_{i}(s,a)$ according to Eq. (11). Specifically, the closer the trajectory is from $\mathcal{M}_{E}$, the larger the weight of $\tau$. Therefore, such a trajectory will contribute more to the smooth guidance reward. A new smoothed RL objective is obtained by inserting Eq. (9) in the standard RL objective:
@@
Given a trajectory with the outcome, how can we determine the relevance of each state-action pair toward achieving this outcome in tasks with sparse or delayed rewards? This is the central topic that temporal credit assignment copes with. Our method can potentially perform long-term credit assignments. Inspired by LfD methods, we propose to realize this goal by using state-only demonstration data in sparse reward settings. Instead of learning complex models to reveal the relevance of action to the future result or adopting contribution analysis methods to achieve return redistribution, our method learns a guidance reward function to assist policy optimization.
Our method can be regarded as a simple redistribution method - the guidance reward for each state-action pair is obtained by calculating the weighted average of the trajectory returns based on the MMD distance between trajectories. In particular, this design of smooth guidance rewards considers the state distribution information of demonstrations and the trajectory returns simultaneously. More importantly, our method does not require obtaining trajectories with sparse rewards before performing credit assignments, which expands the application scope of CA approaches and provides a new idea to CA.